In attempt to restore the economy to its long run equilibrium, the government has decided to cut spending by $20 million (decrease in government spending of $20 million) and increase its transfer payments by $30 million (increase in government transfers of 30 million). If the marginal prosperity to consume 0.67 (2/3), what will be the affect on real GDP of these two actions.
1
Expert's answer
2020-06-09T17:24:36-0400

GDP = Y= C + I + G + (X-M)

MPC = 0.67

Government expenditure/spending

Multiplier effect = 1/1-MPC

Multiplier effect = 1/(1-0.6667)

Multiplier effect = 1/0.3333 = 3

Decrease in G by $20M

Multiplier effect = -$"20\times 3" = -$60 million 

Transfer payments

Transfer payments reduce the effect of a change in real GDP on disposable personal income.

Increase in transfer payments by $30million = MPC \(1-MPC)

KT = MPC/1-MPC

KT = 0.67/1-0.67 = 2 * $30 = $60 million increase in real GDP

Net effect on real GDP =Increase in government transfers of 30 million + decrease in government spending of $20 million

Net increase of $ in real GDP = $ 60 -$60 = $ 0

Thus, the overall net effect on real GDP is 0. Hence, in the long run, the government will be in a position to restore the economy to equilibrium.